OpenStudy (anonymous):

write 5/3+sqrt2 in the form a and b are rational numbers

OpenStudy (anonymous):

in the form a + b sqrt2 where a and b are rational numbers

OpenStudy (dumbcow):

multiply top/bottom by conjugate (3-sqrt2)

OpenStudy (anonymous):

15-5sqrt2 for the top right?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

idk how to multiply the bottom

OpenStudy (dumbcow):

hint \[(a+b)(a-b) = a^{2} -b^{2}\]

OpenStudy (anonymous):

does srt 2 times srt 2 equal to 4?

OpenStudy (anonymous):

because I did that the first time but my final answer was incorrect

OpenStudy (dumbcow):

no that would be 2*2 :) \[\sqrt{2}*\sqrt{2} = \sqrt{4} = 2\]

OpenStudy (anonymous):

oh, I made a silly mistake lol

OpenStudy (anonymous):

that will mean 7 would be the bottom

OpenStudy (dumbcow):

haha yep

OpenStudy (anonymous):

so that will mean 15/7-5sqrt2/7

OpenStudy (dumbcow):

simple as that...and you're done

OpenStudy (anonymous):

sweet haha, i just confused srt 2 times srt 2 as equal to 4 lol